Project

General

Profile

12052013 Minutes
12-5

Mat, Nick, Micheal, Erik, Dom, Gavin
Operations:
-dcm-5-5, 5-4
-Work on cooling
-Control room
-Network - FTS Network issue from Ash River to FNAL
-Watchdogs page at full white out
-Peter turned of the all-steam.
- Turn up the threshold on the no APD parts of the detector. Currently bare FEBs are >50% for the data.
- Current FEB threshold 15-16
- Before 17-21

Scanning session tomorrow - check on a bigger room.

Dom - POT Counting - We see a factor of 2 between numb spill and Steve’s number. Lisa is going to investigate this.

Scanning tutor - Tomorrow